1. Two dice are rolled. Find the probability both are 3.
2. Two dice are rolled. Find probability 3 or 5.
3. Two cars are selected from a production line of 20 cars where 4 are defective. Please find probability both cars are defective.
4. Two cars are selected from a production line of 20 cars where 4 are defective.
Please find probability first car or second car are defective.

Answer by Boreal(15235)

it is (1/6)(1/6)=(1/36) multiply when independent.
3 or 5 is (1/3). Add.
It is (4/20)(3/19)=12/380=3/95.
---------------4/20---------3/19=12/380
|---------------------------16/19=64/380
|--------------16/20--------4/19=64/380
|---------------------------15/19=240/380
The top three are the first or the second (or both) which is 140/380 or 7/19.
If one means either one or not both, then it is 120/380 or 6/19. That has to be clarified in the problem.